Required information

[The following information applies to the questions displayed below.]

During the year, TRC Corporation has the following inventory transactions.

Date Transaction Number of Units Unit Cost Total Cost
Jan. 1 Beginning inventory 60 $ 52 $ 3,120
Apr. 7 Purchase 140 54 7,560
Jul. 16 Purchase 210 57 11,970
Oct. 6 Purchase 120 58 6,960
530 $ 29,610


For the entire year, the company sells 450 units of inventory for $70 each.

3. Using weighted-average cost, calculate ending inventory, cost of goods sold, sales revenue, and gross profit. (Round "Average Cost per unit" to 2 decimal places and all other answers to the nearest whole number.)

Sales revenue
Gross profit
Answer #1

Average cost per unit = 29,610/530 = 55.87 per unit

Cost of Goods Sold = 450 * 55.87 = 25,142

Sales revenue (450*70) 31,500
Gross profit (31,500 - 25,142) 6358

Comment if you face any issues